### v2

# Optimizing Integrated Terrestrial and Non-Terrestrial Networks Performance with Traffic-Aware Resource Management

Abstract
To address an ever-increasing demand for ubiquitous high-speed connectivity, mobile network deployments are becoming increasingly dense. However, this densification has also led to a surge in overall energy consumption, making the process increasingly challenging. In recent years, non-terrestrial networks have been mainly endorsed as a potential solution to enhance coverage by complementing the coverage of the terrestrial network (TN) in areas with limited network deployment. However, their ability to reduce TN energy consumption, though often overlooked, remains a significant advantage. To this end, this paper introduces a novel radio resource management algorithm, BLASTER (Bandwidth SpLit, User ASsociation, and PowEr ContRol), which integrates bandwidth allocation, user equipment (UE) association, power control, and base station activation within an integrated terrestrial and non-terrestrial network (TN-NTN). This algorithm aims to optimize network resource allocation fairness and energy consumption dynamically, demonstrating new opportunities in deploying satellite networks in legacy cellular systems. Our study offers a comprehensive analysis of the integrated network model, emphasizing the effective balance between energy saving and Quality of Service (QoS), and proposing practical solutions to meet the fluctuating traffic demands of cellular networks.
Index Terms:
NTN, LEO Satellites, Load Balancing, Network Energy Efficiency, Resource Allocation.
I Introduction
In recent years, the rapid advancement of cellular communications has driven a marked increase in the demand for high-speed data connectivity. This surge has led to stringent requirements for achieving widespread network connectivity and delivering enhanced network capacity. To meet these challenges, mobile operators have intensified terrestrial macro base stations deployment. However, this approach has its limits, as coverage cannot be guaranteed in logistically challenging locations [1]. In addition, the extensive deployment of TNs with ever-increasing capabilities to deliver high capacity also escalates the overall network energy consumption—a concern in light of current environmental and economic conditions. Consequently, a principal aim in the development and management of mobile networks is minimizing energy consumption, while adhering to QoS standards [2]. Fortunately, recent advancements related to NTNs have offered a promising alternative for extending coverage and enhancing capacity. NTNs utilise airborne vehicles such as unmanned aerial vehicles, high-altitude platform station (HAPS), or satellites, which serve as MBSs or relay nodes to enable connectivity for UEs across the network. The primary benefit of NTNs is their capability to provide broad coverage over large and/or remote areas, where establishing terrestrial MBSs would be either too expensive or difficult. Among the different options for deployment, low-earth orbit (LEO) satellites are poised to become the leading method to achieve high-capacity connectivity from space [3, 4]. In fact, orbiting at altitudes between $200$ and $2000$ kilometers, their relative proximity to Earth provides enhanced signal strength and lower latency compared to other types of constellations. Consequently, this proximity leads to reduced energy needs for launching and lower power usage for transmitting signals to and from the satellite [4]. Spearheaded by companies like SpaceX and OneWeb, ongoing LEO constellation projects aim to launch thousands of LEO satellites around the Earth to establish an ultra-dense constellation. Through a collaboration with mobile operators, the objective is to create an integrated TN-NTN, which can deliver seamless and high-capacity communication services [5], as well as guarantee efficient services for UEs in the future [6, 7].
Therefore, in the context of an integrated TN-NTN, a better understanding of the collaboration between terrestrial and non-terrestrial tiers and the optimization of radio resource management is needed, to fully leverage the perks LEO satellite constellation can offer.
I-A Related works
The resource allocation problem in an integrated TN-NTN has garnered some interest in the last few years. The authors in [8] have introduced a novel radio resource management scheme designed to enhance the performance of an integrated TN-NTN system using 5G 3rd Generation Partnership Project (3GPP) new radio (NR). The proposed solution leverages multicast subgrouping techniques to group UEs within the coverage areas of both terrestrial and non-terrestrial MBSs, ensuring optimal resource allocation to boost overall network performance. The aim was to reduce the UE ping-pong effect between both tiers, increase the resilience to frequent handovers, and simultaneously improve the quality of service perceived by UEs located at cell edges, ultimately enhancing their data rate experience.
Cooperative terrestrial-satellite transmissions are also discussed in [9] where beamforming and frequency reuse are utilised to minimise resource consumption, when terrestrial and satellite MBSs collaboratively deliver services to ground UEs, grouped by their desired content. In contrast, the authors of [10] have explored the joint application of multigroup precoding and resource allocation to create multiple UE groups that are served by either terrestrial or satellite MBSs in different time slots, aiming to reduce interference, while reusing the same frequency.
Indeed, interference management is a critical challenge to achieve efficient resource allocation, which must be addressed to ensure both tiers of the integrated TN-NTN can coexist and achieve optimal performance. To that end, studies such as [11, 12, 13] have provided important perspectives. The authors of [11] have developed a stochastic model to assess interference from terrestrial MBSs to satellite services in the upper $6$ GHz spectrum. The authors have introduced both stochastic and geometrical-stochastic interference models to evaluate the aggregated interference, ensuring it remains within acceptable limits for geostationary satellites. In [12], the authors have investigated the interference between terrestrial and non-terrestrial tiers in the $12$ GHz band, focusing on the potential impact of 5G MBSs on satellite systems. The authors have developed a simulation framework incorporating real-world deployment data and proposed strategies like exclusion zones and beamforming to mitigate interference and enable coexistence. Finally, [13] has presented a framework for spectrum sharing between beyond 5G TNs and satellite systems, optimizing MBS parameters such as transmit power and beam selection based on contextual factors like weather and satellite trajectory. The framework enhances spectrum utilization, while maintaining interference control, outperforming traditional, static spectrum-sharing policies.
Load Balancing in TNs is a topic that has also been well studied over the past few years, but in the integrated TN-NTN scenario, the contributions are limited. Typical methods used for load balancing involve the optimization of a utility function through a pricing-based association strategy [14, 15]. In this line, the authors in [7] have examined an integrated TN-NTN set up in an urban setting, and have shown that diverting some of the traffic to LEO satellites improves the overall signal quality and decreases outages accordingly. The authors of [16] and [17] have exploited the qualities of HAPSs to improve the QoS of ground UEs. [16] has proposed a fairness optimization approach for integrated TN-NTNs, using multiple-input multiple-output (MIMO) beamforming to improve spectral efficiency and manage interference, demonstrating superior performance over standalone TNs. In contrast, [17] has explored multi-connectivity offloading strategies using UAVs and HAPS in NTNs to reduce task computation latency for time-sensitive applications.
In [18] and [19], the authors have investigated the uplink performance of an integrated TN-NTN, leveraging LEO satellites to provide backhaul support to terrestrial MBSs. Both studies aimed to maximise the total uplink data rate, while adhering to backhaul capacity limits. To do this, [19] has taken into account minimum rate requirements, and has adjusted the bandwidth division between fronthaul and backhaul links, while [18] has enhanced UE association and power management using matching algorithms.
In our previous work [20], we have extended the load balancing literature in an integrated TN-NTN. We have introduced a framework using a pricing-based association which, with the assistance of satellites, has managed to successfully distribute the network load, increase the maximum network throughput, and enhance the network coverage.
As stated previously, in the current environmental context, improving network energy efficiency (EE) and reducing energy consumption have become major objectives. To this end, mobile operators need to adapt the offered capacity by the terrestrial MBSs to the rate requirements by dynamically adjusting the number of active MBSs. For example, operating all MBSs during low-traffic periods is suboptimal, as keeping underused or idle terrestrial MBSs active leads to wasteful use of energy and communication resources [2]. In the context of an integrated TN-NTN, it is also beneficial to deactivate some terrestrial MBSs and handover the UEs to satellites to decrease the energy consumption of the TN.
MBSs activation is a well-studied topic in TNs. The authors of [21] have introduced an energy-efficient algorithm that strategically shuts down MBSs one at a time, ensuring they do not overburden neighbouring MBSs. To preserve the QoS, [22] has examined the effects of traffic offloading in heterogeneous networks on energy use, and proposed a centralized Q-learning method to strike a balance between energy saving and QoS satisfaction. In addition, the authors of [23] have devised an algorithm that enables UEs to associate with multiple MBSs across different frequency bands, simultaneously optimising the transmit power of the MBSs to facilitate their shutdown during periods of low traffic. The authors of [24] have also tackled traffic uncertainties in ultra-dense networks by optimising both MBS activation and UE association strategies, employing chance constraint programming based on statistical traffic data to effectively balance traffic loads and reduce interference. Recent studies such as [25, 26, 27] have explored the integration of HAPS to enhance network efficiency in a more dynamic and sustainable manner. Indeed, [25] has investigated how HAPSs can complement traditional network densification to manage dynamic traffic in urban areas, demonstrating better energy efficiency and sustainability by using HAPSs to handle peak demand periods. In [26], the authors have tackled the traffic load estimation issue in HAPS-assisted networks, proposing Q-learning algorithms to optimize cell-switching strategies, improving energy efficiency and making advanced cell-switching methods feasible for vertical heterogeneous networks. The authors of [27] have investigated the challenges of MBSs activation in an integrated TN-NTN using HAPS. The authors have focused on offloading traffic from deactivated terrestrial MBSs to the HAPS, mainly using a sorting algorithm, which prioritises switching off MBSs with relatively lower traffic loads. Although the studies in [25, 26, 27] have provided valuable insights and promising results, they have not explored the optimization of spectrum sharing and allocation strategies, thereby limiting resource utilization efficiency in integrated TN-NTNs. In our latest work [28], we have introduced a framework which leverages the large coverage ability of LEO satellites to shutdown terrestrial MBSs in rural areas during low traffic. To the best of our knowledge, only the authors of [26], [27], and [28] have considered NTNs as a solution for meeting coverage and capacity demands while deactivating MBSs.
I-B Contributions
In this paper, we consider an integrated TN-NTN, and introduce a novel radio resource management algorithm that adapts to the fluctuating traffic of the network. The contributions of this paper are summarized as follows:
  • We develop BLASTER (Bandwidth SpLit, User ASsociaTion, and PowEr ContRol), an innovative approach to adaptive radio resource management. BLASTER seamlessly integrates the management of bandwidth allocation and UE association between terrestrial and non-terrestrial tiers. It also controls the transmission power and activation of terrestrial MBSs. The proposed framework is designed to balance network fairness and energy consumption by adjusting to current traffic conditions. Our findings demonstrate that BLASTER can reduce the overall TN energy usage by $67\,\%$ when compared to a standard integrated TN-NTN system adhering to 3GPP guidelines, while notably enhancing the average network sum log-throughput (SLT) in times of high traffic demand by at least $6\,\%$ compared to the same 3GPP benchmark mentioned above.
  • We propose a framework where the terrestrial and non-terrestrial tiers orthogonally share the same frequency band based on the fluctuating traffic. In this scenario, we demonstrate that the optimal bandwidth allocation for the non-terrestrial tier is proportional to the fraction of UEs associated to the satellites, provided that UEs have the same requirements. Through dynamic allocation of the resources for both tiers, we are able to optimize the collaboration of both tiers and enhance the network efficiency throughout the day.
  • By exploiting the special properties of the formulated optimization problem, we design a practical heuristic with an intuitive behaviour to solve the problem with limited complexity, achieving results that highlight the trade-off between enhancing network SLT and reducing energy consumption.
Note that, since recent techno-economic [29, 30] studies have underscored the cost-effectiveness of the LEO satellites compared to HAPS, which are still suffering from technological, regulatory and economic constraints, we have focused on LEO satellites in this paper, but BLASTER is also applicable to HAPSs.
The remainder of the paper is as follows: Section II introduces the system model used in this work. In Section III, we formulate the problem that we aim to solve, and detail the proposed solutions in Section IV. In Section V, we study the performance of both developed frameworks and compare them to standard benchmarks. The paper is concluded in Section VI.
II System Model
Our study focuses on the downlink (DL) of a cellular network which consists of $M$ terrestrial MBSs and $N$ MBSs installed on a constellation of LEO satellites, making a total of $L$ MBSs. They provide service to $K$ UEs located in the area of study, which is composed of a rural and urban zone. We refer to the overall network bandwidth as $W$, which is distributed by the mobile network operator across the ground and space-based network. We suppose that the network operates within the S band, around 2 GHz, with ground and satellite MBSs using orthogonal, but dynamically adjustable, portions of this band. Throughout this paper, we will use $\mathcal{T}$ and $\mathcal{S}$ to represent the set of terrestrial and satellite MBSs respectively. Moreover, $\mathcal{B}=\mathcal{T}\cup\mathcal{S}=\{1,\cdots,j,\cdots,L\}$ is the complete set of MBSs, and $\mathcal{U}=\{1,\cdots,i,\cdots,K\}$ defines the set of UEs. For the remainder of the paper, we will designate the Hadamard product with $\odot$.
II-A Channel Model
We model the channel for both terrestrial and satellite links based on the 3GPP recommendations provided in [31, 32].
II-A1 Terrestrial link channel model
The large-scale channel gain between a ground-based MBS $j$ and a UE $i$ is calculated as follows:
$\begin{split}\beta_{ij}&=\lfloor G_{T_{X}}+G_{UE}+PL_{ij}^{b}+SF_{ij}+PL_{ij}^{tw}+PL_{ij}^{in}\\ &+\mathcal{N}\left(0,\sigma^{2}_{p}\right)\rceil,\end{split}$ (1)
[…]
II-A2 Satellite link channel model
Similarly, if a satellite MBS $j$ is serving a UE $i$, we can compute the large-scale channel gain (detailed in [32]) as:
$\beta_{ij}=\left\lfloor G_{T_{X}}+G_{UE}+PL_{ij}^{b}+SF_{ij}+CL+PL_{ij}^{s}+PL_{ij}^{e}\right\rceil.$ (2)
In (2), $CL$ accounts for clutter loss, which is the attenuation arising from obstacles such as buildings and vegetation surrounding the UE. $PL_{ij}^{s}$ captures the scintillation loss, reflecting the quick changes in signal amplitude and phase due to ionospheric disturbances. Finally, $PL_{ij}^{e}$ refers to the building entry loss, an attenuation that occurs for all UEs located indoors.
It should be noted that components $SF_{ij}$ and $CL$ are function of -and greatly vary with- the LoS conditions, and that the LoS probability for the satellite links are provided in [32, Table 6.6.1-1]. Additionally, it should be highlighted that the elevation angle of the satellite also impacts the quality of the channel link. The elevation angle is the angle between the horizontal plane (the plane parallel to the surface of the Earth at the UE location) and the line of sight to the satellite. If the elevation angle increases, $CL$, $PL_{ij}^{e}$ and $SF_{ij}$ are greatly reduced. Given the cartesian coordinates of a satellite $s$ $\left(x_{s},y_{s},z_{s}\right)$ and a UE $u$ $\left(x_{u},y_{u},z_{u}\right)$, we can compute the elevation angle $\theta_{u}$ as:
$\theta_{u}=\arcsin\left(\frac{z_{s}-z_{u}}{\sqrt{\left(x_{s}-x_{u}\right)^{2}+\left(y_{s}-y_{u}\right)^{2}+\left(z_{s}-z_{u}\right)^{2}}}\right).$ (3)
[…]
Considering that each UE is either associated to a terrestrial or satellite MBS, and there is no interference between the two tiers, as they are allocated different bandwidths orthogonally, we can calculate the large-scale signal-to-interference-plus-noise ratio (SINR) for each UE $i$ and MBS $j$ as:
$\gamma_{ij}=\frac{\beta_{ij}p_{j}}{\sum\limits_{\begin{subarray}{c}j^{^{\prime}}\in\mathcal{I}_{j}\end{subarray}}\beta_{ij^{^{\prime}}}p_{j^{^{\prime}}}+\sigma^{2}}.$ (4)
In (4), $p_{j}$ denotes the transmit power allocated per resource element (RE) at MBS $j$, $\mathcal{I}_{j}$ indicates the set of MBSs interfering with the serving MBS $j$ and $\sigma^{2}$ represents the noise power per RE. Moreover, assuming that MBS $j$ equally shares its total available bandwidth $W_{j}$ between the $k_{j}$ UEs it is serving, we are able to compute the mean throughput for UE $i$ connected to MBS $j$ as follows:
$R_{ij}=\frac{W_{j}}{k_{j}}\log_{2}(1+\gamma_{ij}).$ (5)
II-C Energy Consumption Model
While [33] provided one of the most widely used models for 4G MBS energy consumption, it is not well suited for 5G MBSs integrating massive MIMO technology. Consequently, we chose the more recent model proposed in [34] which accounts for massive MIMO and carrier shutdown, and thus fits better to our system design. Note that, in the context of our paper, shutdown and sleep mode are considered equivalent, as both refer to states in which a MBS significantly reduces or halts its operations to conserve energy. More generally, it only takes $3$ seconds to shutdown or wake up a MBS, while the shutdown duration may range from tens of seconds to minutes or even hours [35].
The energy consumption of a MBS can be modelled as the sum of multiple components. We denote as the baseline energy consumption, the energy used by the components that are kept active in a shutdown MBS. Then, we denote as the static component, the energy consumption that occurs regardless of the level of the MBS traffic load. The static energy consumption represents the minimum power required to keep essential systems operational and maintain standby readiness. Finally, the dynamic component refers to the load-dependent energy consumption that fluctuates depending on the MBS traffic load. Typically, the dynamic component increases whenever a MBS increases its transmit power or uses additional transmission resources, e.g., more resource blocks. For a MBS $j$, this model can be formulated as:
$Q_{j}(p_{j})=P_{0}+p_{j}+\psi_{j}||p_{j}||_{0},$ (6)
where $P_{0}$ represents the baseline energy consumption, $\psi_{j}$ represents the static component and $p_{j}$ accounts for the dynamic consumption of the MBS. Also, $||\cdot||_{0}$ is a binary-valued function equal to $1$ if the transmit power $p_{j}$ is greater than 0.
Regarding the LEO satellite, the total energy consumption can be expressed as the sum of the inherent energy consumption of the LEO satellite, which accounts for altitude adjustments, GPS navigation, and routing operations, plus the energy consumption of the MBS installed on it, as detailed in (6). We suppose that the satellites in the constellation are solar-powered and well-dimensioned. Thus, they can handle both the power requirements needed for an operational satellite and manage the telecom equipment added as payload, based on the growing adoption of real-world projects such as Starlink or Kuiper. The major notations are summarized in Table I.
TABLE I: List of Notations
Parameter Symbol
Carrier frequency $f_{c}$
Total number of MBSs $L$
Total number of UEs $K$
Carrier bandwidth $W$
Set of all MBSs $\mathcal{B}$
Set of all UEs $\mathcal{U}$
Large-scale channel gain for MBS $j$ and UE $i$ $\beta_{ij}$
Elevation angle of the satellite relative to the UE $\theta_{u}$
Noise power per RE [dBm] $\sigma^{2}$
SINR for MBS $j$ and UE $i$ $\gamma_{ij}$
Mean throughput for UE $i$ perceived from MBS $j$ $R_{ij}$
Perceived throughput for UE $i$ $R_{i}$
Energy consumption for MBS $j$ $Q_{j}$
Transmit power per RE at MBS $j$ $p_{j}$
Maximum transmit power per RE at MBS $j$ ${p_{j}}_{\mathrm{max}}$
Vector with transmit power per RE of every MBS $p\in\mathbb{R}^{L}$
Binary matrix for UE-MBS association $X\in\mathbb{R}^{K\times L}$
Fraction of bandwidth allocated to satellite MBSs $\varepsilon$
Coverage Threshold [dBm] $RSRP_{\text{min}}$
Regularization parameter $\lambda$
Hadamard product $\odot$
III Problem Formulation
Our objective is to develop a framework, which simultaneously increases UE performance and reduces network energy consumption by adjusting the resource distribution between satellite and terrestrial MBSs in response to the hourly fluctuations in network traffic. Specifically, our goal is to achieve proportional fair resource allocation by optimizing the network SLT, while limiting the TN energy consumption: in fact, the nature of the logarithmic cost function discourages each MBS to allocate disproportionate resources to a single UE. Let us denote as $\varepsilon$ the fraction of the bandwidth allocated to the LEO satellites at a given hour of the day; then, the bandwidth allocated for an MBS $j$ can be written as:
$\begin{split}W_{j}=\begin{cases}\varepsilon W&\text{if}\ j\in\mathcal{S},\\ \left(1-\varepsilon\right)W&\text{else}.\end{cases}\end{split}$ (7)
Also, we define $x_{ij}$, a binary variable, which equals $1$ if UE $i$ is associated to MBS $j$. Accordingly, the perceived throughput for UE $i$ can then be reformulated as:
$R_{i}=\sum_{j\in\mathcal{B}}x_{ij}R_{ij.}$ (8)
[…]
IV Designed Solution
In this section, we first present BLASTER, the framework proposed to address the optimization problem outlined in (III)-(III). Then, we design a low-complexity heuristic based on the special characteristics of the problem, and provide a comparison for BLASTER w.r.t. performance state-of-the-art benchmarks. Finally, we provide an analysis of the complexity of both solutions.
IV-A BLASTER
We adopt the block coordinate gradient ascent (BCGA) algorithm to solve problem (III)-(III). BCGA is a technique used to maximize a function by iteratively updating its different parameters. In our case, we begin by optimizing the UE-MBS association and bandwidth split while considering the transmit power fixed. Then, we optimize the transmit power at each MBS whilst keeping the two previous parameters unchanged. The full overview is provided in Algorithm 1.
Data: K UEs and L MBs.
Initialization: s = 00;
X: Association done through max-RSRP;
p: Transmit power set to maximum;
$\varepsilon=0.5$;
  // Equal bandwidth split
Compute: $f\left(X,\varepsilon,p\right)$ // Initial point
$w=\left[1,\dots,1\right]\in\mathbb{R}^{L}$;
Initialize $\alpha\in\mathbb{R}^{K\times L}$, $\mu\in\mathbb{R}^{K}$;
Initialize $\eta\in\mathbb{R}^{L}$;
Initialize $\delta\in\mathbb{R}$;
while Utility function $f$ has not converged do
       // UE Association and bandwidth split
       Compute: $\tilde{X}(s)=X(s)+\alpha\nabla_{X}f\left(X,p,\varepsilon\right)$   (9);
       Solve (IV-A1) using gradient projection to obtain $\mu^{*}$;
       Compute:
       $=\max\{\tilde{X}(s)-\beta\odot p^{\rm PAD}\odot{\mu^{*}}^{\rm PAD},0\}(\ref{X_s+1})$ ( );
       $\varepsilon^{*}=\frac{K_{\mathcal{S}}}{K}$                                                         (20);
       // Power control step
       Compute: $\tilde{p}(s)=p(s)+\eta\nabla_{p}f\left(X,p,\varepsilon\right)\quad$ (21);
       Compute: $t=\lambda\cdot\eta\cdot w^{T}\psi\quad$                          (22);
       Compute: $\hat{p}(s)=\rm\max\bigl{\{}1-\frac{t}{||\tilde{p}(s)||_{2}},0\bigr{\}}\tilde{p}(s)$    (23);
       Compute: $\tau$ based on (25);
       Compute: $p{(s+1)}=\Bigr{[}\hat{p}(s)\Bigr{]}_{\tau}^{p_{\rm max}}$                   (26);
       $w=\left[\frac{1}{p_{1}+\delta},\dots,\frac{1}{p_{L}+\delta}\right]$;
       // $\delta$ small constant to avoid numerical instability
       Compute: $f\left(X(s),\varepsilon,p(s)\right)$ ;
       $s=s+1$;
end while
return X,$\varepsilon$,p;
Algorithm 1 BLASTER Framework
IV-A1 Utility optimization under fixed transmit power
Let $f$ represent the utility function we aim to maximize in (III). By relaxing constraint (III) such that $x_{ij}\in\left[0,1\right]$, we get a convex optimization problem with respect to $X$. To tackle this problem, the iterative gradient projection method serves as an ideal solution [37], given its suitability for constrained optimization. The gradient projection method involves calculating the gradient of the objective function and then projecting this gradient onto the viable region delineated by the problem constraints. Adopting the gradient projection method, we can determine the gradient update at iteration $s$ as follows:
$\tilde{X}(s)=X(s)+\alpha\nabla_{X}f\left(X,p,\varepsilon\right),$ (9)
[…]
$\begin{split}\mathcal{L}\left(X,\mu\right)&=\frac{1}{2}||X-\tilde{X}||^{2}_{F}+\left(\tilde{\beta}\cdot p-RSRP_{\rm min}\cdot\mathbbm{1}_{K}\right)^{T}\mu\\ &=\frac{1}{2}||X||^{2}_{F}-\mathrm{Tr}\left(X^{T}\tilde{X}\right)+\frac{1}{2}||\tilde{X}||^{2}_{F}+\left(\tilde{\beta}\cdot p\right)^{T}\mu\\ &-\left(RSRP_{\mathrm{min}}\cdot\mathbbm{1}_{K}\right)^{T}\mu,\end{split}$ (10)
where $\mu\in\mathbb{R}^{K}$ is the Lagrange multiplier associated with constraint (III). Calculating the gradient of (10) with respect to $X$, we obtain:
$\begin{split}\nabla_{X}\mathcal{L}\left(X,\mu\right)=X-\tilde{X}+\beta\odot\underbrace{\left(\mathbbm{1}_{K}\cdot p^{T}\right)}_{:=p^{\rm PAD}}\odot\underbrace{\left(\mu\cdot\mathbbm{1}_{L}^{T}\right)}_{:=\mu^{\rm PAD}}.\end{split}$ (11)
Fixing the dual variable, we are able to determine the minimal point for the gradient, which is obtained for:
$\begin{split}X^{\star}=\max\{\tilde{X}-\beta\odot p^{\rm PAD}\odot\mu^{\rm PAD},0\}.\end{split}$ (12)
Subsequently, we can introduce the Lagrangian dual function, formulated as:
$\begin{split}&\mathcal{D}\left(\mu\right)=\max\limits_{X}\mathcal{L}\left(X,\mu\right).\end{split}$ (13)
Proposition 1.
We can rewrite $\mathcal{D}\left(\mu\right)$ as:
$\begin{split}\mathcal{D}\left(\mu\right)&=\frac{1}{2}||X^{\star}||^{2}_{F}-\mathrm{Tr}\left(X^{\star}\left[\tilde{X}-\beta\odot p^{\rm PAD}\odot\mu^{\rm PAD}\right]^{T}\right)\\ &-\left(RSRP_{\mathrm{min}}\cdot\mathbbm{1}_{K}\right)^{T}\mu.\end{split}$ (14)
Proof.
Please refer to Appendix A. ∎
Also, we notice that for any matrix $A$:
$\frac{1}{2}\Big{|}\Big{|}\max\{A,0\}\Big{|}\Big{|}^{2}_{F}-\mathrm{Tr}\left(\max\{A,0\}A^{T}\right)=-\frac{1}{2}\Big{|}\Big{|}\max\{A,0\}\Big{|}\Big{|}^{2}_{F}$ (15)
[…]
$X(s+1)\triangleq X^{\star}=\max\{\tilde{X}(s)-\beta\odot p^{\rm PAD}\odot{\mu^{*}}^{\rm PAD},0\}.$ (16)
We repeat those iterations until convergence to obtain the optimal association setting $X^{*}$.
Once this is done, we have to optimally split the bandwidth between both terrestrial and non-terrestrial tiers. To this aim, we rewrite the mean UE throughput, defined in (5), as follows:
$\begin{split}R_{ij}=\begin{cases}\varepsilon r_{ij}&\text{if}\ j\in\mathcal{S},\\ \left(1-\varepsilon\right)r_{ij}&\text{otherwise},\end{cases}\end{split}$ (17)
where
$r_{ij}=\frac{W}{k_{j}}\log_{2}\left(1+\gamma_{ij}\right).$
Note that, since in (17) both the MBS transmit power and the noise power scale linearly with the bandwidth, $\gamma_{ij}$ is unaffected by the bandwidth split between satellite and TN.
Proposition 2.
The optimal bandwidth allocation for the non-terrestrial tier linearly increases with the fraction of UEs associated to a satellite MBS in the network, provided that all UEs have the same requirements, i.e.:
$\varepsilon^{*}=\frac{K_{\mathcal{S}}}{K}$ (18)
where $K_{\mathcal{S}}$ denotes the number of UEs associated to a satellite within the network.
Proof.
First, we compute the gradient of our utility function $f$ with respect to $\varepsilon$:
$\begin{split}&\nabla_{\varepsilon}f\left(X,p,\varepsilon\right)=\frac{\partial}{\partial\varepsilon}\left(\sum_{i=1}^{K}\log\left(R_{i}\right)\right)=\sum_{i=1}^{K}\frac{\partial}{\partial\varepsilon}\log\left(R_{i}\right)\\ &=\sum_{i=1}^{K}\frac{\frac{\partial}{\partial\varepsilon}\left(R_{i}\right)}{R_{i}}=\sum_{i=1}^{K}\frac{\frac{\partial}{\partial\varepsilon}\left[\sum\limits_{j\in\mathcal{S}}\varepsilon x_{ij}r_{ij}+\sum\limits_{j\in\mathcal{T}}\left(1-\varepsilon\right)x_{ij}r_{ij}\right]}{R_{i}}\\ &=\sum_{i=1}^{K}\frac{\sum\limits_{j\in\mathcal{S}}\frac{\partial}{\partial\varepsilon}\left[\varepsilon x_{ij}r_{ij}\right]+\sum\limits_{j\in\mathcal{T}}\frac{\partial}{\partial\varepsilon}\left[\left(1-\varepsilon\right)x_{ij}r_{ij}\right]}{R_{i}}\\ &=\sum_{i=1}^{K}\left[\frac{\sum\limits_{j\in\mathcal{S}}x_{ij}r_{ij}-\sum\limits_{j\in\mathcal{T}}x_{ij}r_{ij}}{R_{i}}\right].\end{split}$ (19)
Let us denote as $\mathcal{U_{S}}$ and $\mathcal{U_{T}}$ the sets of UEs served by the satellites and terrestrial MBSs respectively, such that $\mathcal{U}=\mathcal{U_{S}}\ \cup\ \mathcal{U_{T}}$; then, by setting (19) to 00, we can strike the optimal splitting point for the bandwidth as follows:
$\begin{split}&\nabla_{\varepsilon}f\left(X,p,\varepsilon\right)=0\Leftrightarrow\sum_{i\in\mathcal{U_{S}}}\frac{1}{\varepsilon}+\sum_{i\in\mathcal{U_{T}}}\frac{-1}{1-\varepsilon}=0\\ &\Leftrightarrow\frac{K_{\mathcal{S}}}{\varepsilon}-\frac{K-K_{\mathcal{S}}}{1-\varepsilon}=0\Leftrightarrow\varepsilon^{*}=\frac{K_{\mathcal{S}}}{K}.\end{split}$ (20)
IV-A2 Transmit power optimization under fixed association
[…]
$\tilde{p}(s)=p(s)+\eta\nabla_{p}f\left(X,p,\varepsilon\right)$ (21)
where $\eta\in\mathbb{R}^{L}$ is a pertinent step-size. Then, as outlined in [38], the proximal gradient method updates $p$ by addressing the problem described below: {mini!}—s—[2] p12 ——~p(s) - p(s) ——^2_2 + t ——p(s) ——_2 with
$t=\lambda\cdot\eta\cdot w^{T}\psi.$ (22)
A closed-form solution to this problem, known as block-soft thresholding [38, Sec. 6.5.1] is expressed as:
$\hat{p}(s)=\rm\max\bigl{\{}1-\frac{t}{||\tilde{p}(s)||_{2}},0\bigr{\}}\tilde{p}(s).$ (23)
After updating the transmit power vector, it is necessary to project it into a feasible region to ensure compliance with constraints (III) and (III). The maximum transmit power per RE naturally sets the upper limit of our feasible region. To determine the lower boundary, we apply the minimal coverage constraint. Indeed, as indicated by (III), each UE connected to a MBS $j$ must receive a signal power that exceeds the minimum required RSRP value, $RSRP_{\rm min}$. This can be rewritten as:
$\begin{split}&\forall i\in\mathcal{U}_{j},\quad p_{j}\quad{\geq}\quad\frac{RSRP_{\rm min}}{\beta_{ij}},\end{split}$ (24)
with $\mathcal{U}_{j}$ representing the set of UEs connected to the MBS $j$. Consequently, we can define the lower bound of the feasible region for each MBS $j$ as:
$\tau_{j}=\max_{i\in\mathcal{U}_{j}}\left(\frac{RSRP_{\min}}{\beta_{ij}}\right).$ (25)
Finally, the transmit power update at the iteration $s$ can be computed as follows:
$p{(s+1)}=\Bigr{[}\hat{p}(s)\Bigr{]}_{\tau_{j}}^{p_{\rm max}}.$ (26)
After the algorithm yields $p(s+1)$, we adjust the power weights $w_{j}$ following the re-weighting algorithm described in [23]. Indeed, we update the vector $w$ as follows:
$w(s+1)=\left[\frac{1}{p_{1}+\delta},\dots,\frac{1}{p_{L}+\delta}\right]\;$ (27)
where $\delta$ is a parameter introduced to avoid numerical instability. By updating the vector $w$ based on (27), we reduce the impact on the utility function of the MBSs that have a large transmit power to provide continuous coverage to the associated UEs. In contrast, this weighing method effectively pushes MBSs that are not providing traffic to decrease the transmit power further until they shutdown.
IV-B Heuristic development
Given that BLASTER is composed of two iterative methods (BCGA and gradient projection method to solve (IV-A1)), we have developed a low computational complexity heuristic based on the domain expertise. The designed heuristic has a structure similar to that of BLASTER. Indeed, it first deals with the UE-MBS association phase and allocates the bandwidth resources accordingly. Then, it deactivates part of the MBSs in the network and/or updates their transmit power. The heuristic algorithm iteratively repeats these three steps until the selected utility function, i.e., the network SLT converges. Specifically, the three steps of the proposed low computational complexity heuristic can be described as the following:
  • UE Association: To maximize the number of terrestrial MBSs that can be switched off during low-traffic hours (00 AM - $7$ AM), it is desirable to maximize the share of the traffic served by the non-terrestrial tier. Hence, the proposed heuristic associates each UE that has an RSRP larger than $RSRP_{\min}$ to the satellite. For each remaining UEs that is yet to be associated to a MBS, we collect the perceived RSRPs from all the MBSs that are providing an RSRP greater than $\geq RSRP_{\rm min}$. Then, we average these RSRP measurements and rank each UE based on the calculated average from worst to best. Then, starting from the UE with the poorest average, we associate it with the MBS providing the largest throughput with the aim of maximizing the SLT.
    In the case of a high-traffic scenario, the satellite merely acts as an umbrella over the ground network to provide service for UEs that have no service, and potentially facilitate the distribution of the load. In this regard, the UE association process is quite similar to the low-traffic scenario as we rank each UE based on their average perceived RSRP and associate to the MBS providing the best throughput [^2] Note that the association is done only once for each UE and does not change even if the associated MBS is not providing the best throughput anymore due to the increased number of served UEs..
  • Bandwidth split: The bandwidth is split according to the expression derived in (20).
  • MBS shutdown and power control: In low-traffic scenario, the proposed heuristic shuts down all MBSs serving less UEs than a set threshold $T_{UE}$, provided that these UEs can be successfully handed over to neighbouring MBSs. We also reduce the MBS transmit power as long as the coverage constraint (III) is satisfied for the connected UEs (feasible region computed in (25)). On the other side, for high traffic, we only shut down inactive MBSs and reduce the transmit power similarly to the low-traffic scenario.
The complete heuristic is detailed in Algorithm 2.
Data: K UEs and L MBs.
Initialization;
s = 00;
X: Association done through max-RSRP;
p: Transmit power set to maximum;
$\varepsilon=0.5$;
  // Equal bandwidth split
Compute: $SLT$ // Initial point
while $SLT$ has not converged do
       // UE Association and bandwidth split
       if low-traffic hour: then
             for all UEs $u$ do
                   if RSRP perceived from satellite $\geq RSRP_{\rm min}$ then
                         Associate UE $u$ to the satellite;
                   end if
             end for
       end if
      Rank each UE according to the average perceived RSRP;
       for all UEs $u$ uncovered do
             $\mathcal{C}$ : List of MBSs providing an RSRP greater than $RSRP_{\rm min}$ for UE $u$ ;
             Associate $u$ to the MBS $\in\mathcal{C}$ providing the largest throughput.
       end for
      $\varepsilon^{*}=\frac{K_{\mathcal{S}}}{K}$                                                         (20);
       // Power control step
       if low-traffic hour: then
             for all MBSs $b$ serving less than $T_{\mathrm{UE}}$ do
                   if we can handover every UE served by $b$ then
                         Offload each UE to a neighboring MBS;
                         Shutdown MBS $b$;
                   end if
             end for
       end if
      // For remaining active base stations:
       Compute: $\tau$ based on (25);
       // Reduce Tx Power while ensuring coverage for all UEs:
       Compute: $p{(s+1)}=\tau$;
       $s=s+1$;
end while
Return X,$\varepsilon$,p;
Algorithm 2 Heuristic
IV-C Complexity Analysis
Concerning the algorithmic complexity, we can compare both solutions that we have developed:
IV-C1 BLASTER complexity
First, let’s analyse the complexity of the operations in an iteration of BLASTER. For the optimization of $X$, the computation of the gradient step (9) has a complexity of $\mathcal{O}\left(K\times L\right)$. In addition, we solve a gradient projection problem (IV-A1) at the end of the UE association phase using an iterative method. Each iteration of the gradient projection method is executed with a complexity of $\mathcal{O}\left(K\times L\right)$. We denote as $I_{\mu}$ the number of iterations needed to reach a stopping criterion for the gradient projection method. The stopping criterion is either the convergence of the utility function (SLT) or the completion of a given number of iterations. Then, the overall complexity for the optimization of $X$ is $\mathcal{O}\left(K\times L+I_{\mu}\times K\times L\right)=\mathcal{O}\left(I_{\mu}\times K\times L\right)$. To obtain the optimal split of the bandwidth as per (20), our algorithm computes the fraction of the UEs associated to the satellites, which is an operation of complexity $\mathcal{O}\left(K\right)$. Finally, the transmit power vector optimization has an overall complexity of $\mathcal{O}\left(K\times L\right)$. Taking all of this into account, if $I_{\rm blaster}$ represents the number of iterations needed to meet a stopping criterion for BLASTER (specified above), the complexity of the framework is $\mathcal{O}\left(I_{\rm blaster}\times I_{\mu}\times K\times L\right)$.
IV-C2 Heuristic complexity
For modelling the heuristic complexity, we consider separately low and high traffic cases. In the low-traffic scenario, the proposed algorithm associates to a satellite each UE perceiving an RSRP greater than $\geq RSRP_{\rm min}$ from a satellite. This results in having to do the ranking and MBS association process detailed above for $K-K_{\mathcal{S}}$ UEs. Therefore, in low traffic, the complexity of the UE association phase is $\mathcal{O}\left((K-K_{\mathcal{S}})\times L\right)$. Also, the complexity of the operations for the transmit power optimization and BS activation is $\mathcal{O}\left(K\times L\right)$. Therefore, the overall complexity for one iteration in a low-traffic scenario is $\mathcal{O}\left(K\times L\right)$. For the high-traffic case, the complexity for those steps is $\mathcal{O}\left(K\times L\right)$, explained by the fact that the UE association process is done for each of the UEs. Finally, denoting by $I_{H}$ the number of iterations needed for the utility function to converge, the overall complexity of the heuristic is $\mathcal{O}\left(I_{H}\times K\times L\right)$.
V Simulation Results & Analysis
In this section, we assess the performance of the proposed solution under different traffic conditions. Specifically, in Sec. V-A, we describe the simulation settings and the baseline algorithms, and in Sec. V-B we discuss the impact of the regularization parameter $\lambda$, introduced in (III), on the targeted performance metrics. In Sec. V-D, we describe how the algorithms under investigation leverage the satellite network resources to offload the TN. Finally, in Secs. V-E and V-F, we analyse the performance of the solutions under tests in terms of network sum-log throughput and TN energy consumption, respectively.
V-A Simulation settings and benchmarks
All the results and analyses presented in the following section have been conducted over a $24$-hour observation window, with snapshots of the network taken at the start of each hour, and generated using a custom-built system-level simulator, which adheres to 3GPP recommendations [39, 32, 40, 31, 41, 42]. The area of study is approximately $2500$ $\mathrm{km}^{2}$, which matches the beam diameter of a LEO satellite [40], and includes both an urban and a rural area, as depicted in Fig. 1. In this area, we model the traffic variations by changing the number of active UEs in the network. Precisely, the number of UEs deployed in the network scales according to the daily downlink traffic load pattern presented in [43]. The number of UEs decreases during the night, until it reaches its minimum of $400$ at $5$ AM, and increases in high traffic, reaching up to $10000$ at $8$ PM. The UEs are deployed uniformly across the entire study area, with a higher density of deployment in the urban area relatively to the rural one. The terrestrial MBSs in both urban and rural areas are deployed in a hexagonal grid layout [44], with a higher density of MBSs in the urban area. In our simulation setting, we consider that $80\%$ of the UEs are indoor [31].
Figure 1: Representation of area of study.
V-A1 Satellite deployment
In our study, we consider a LEO satellite constellation employing earth-fixed beams [32, Section 4.6], implying that the satellites use beam pointing mechanisms (mechanical or electronic steering feature) to compensate for their mobility and cover a given fixed region. Specifically, we assume in our study that a satellite has a seven beam configuration, as depicted in Fig. 2, and that at any given time, a satellite in the constellation is able to serve the region of interest in red.
As stated in Section II, the elevation angle of a satellite is an important parameter, which impacts the quality of the signal perceived from the satellite. Taking this into account, we embrace the mobility of the LEO satellites by assessing network performance at three distinct positions (P1, P2, P3), as also illustrated in Fig. 2:
  1. P1)
    We make the first performance assessment when the satellite is located at the nadir point of the beam adjacent to our area of study, which is located at a horizontal distance of $50$ kms (as well as a total distance of approximately $603$ kms) from the center of the area of study, with an elevation angle (for a UE located at the center of the central beam) approximately equal to $84^{\circ}$.
  2. P2)
    The second assessment is made when the satellite is at the nadir point, exactly above the center of our area of study.
  3. P3)
    Similarly to the first measurement, we make the assessment when the satellite is $50$ kms away from the center of the area of study, at the nadir point of the next contiguous beam.
Figure 2: Representation of the satellite positions considered for the area of study (in red).
Averaging the values obtained from these three assessments, we can get an overview of the satellite channel quality.
Figure 3: Sum Log Throughput distribution achieved by BLASTER in high traffic for various $\lambda_{\rm max}$.
Figure 4: Sum Log Throughput distribution achieved by BLASTER in low traffic for various $\lambda_{\rm max}$.
V-A2 Benchmarks
To compare and assess the performance of both developed algorithms, we introduce the two following benchmarks: The 3GPP-TN configuration includes only a TN operating on a bandwidth of $10$ MHz. In contrast, the 3GPP-NTN configuration includes a satellite network overlaying the terrestrial one. In this case, the bandwidth is divided according to the 3GPP suggestions in [40], with $30$ MHz assigned to the satellite tier and $10$ MHz to the terrestrial one. In both configurations, the UEs associate to the MBSs based on the max-RSRP rule, with no DL transmit power optimization or MBS shutdown implemented. The most important simulation parameters are listed in Table II following [39, 32, 40, 31, 41, 42].
Parameter Value
Total Bandwidth $W$ $40$ $\mathrm{MHz}$
Carrier frequency $f_{c}$ $2$ $\mathrm{GHz}$
Subcarrier Spacing $15$ $\mathrm{kHz}$
Urban/Rural Inter-Site Distance $500/1732$ $m$
Number of Macro BSs $1776$
Satellite Altitude [40] $600$ km
Terrestrial Max Tx Power per RE ${p}_{\mathrm{max}}$ [41] 17.7 $17.7\text{ }\mathrm{dBm}$
Satellite Max Tx Power per RE ${p}_{\mathrm{max}}$ [40] 15.8 $15.8\text{ }\mathrm{dBm}$
Antenna gain (Terrestrial) $G_{T_{X}}$ [42] 14 $14\text{ }\mathrm{dBi}$
Antenna gain (Satellite) $G_{T_{X}}$ [40] 30 $30\text{ }\mathrm{dBi}$
Shadowing Loss (Terrestrial) $SF$ [31] 4  8 $4\text{ }-\text{ }8\text{ }\mathrm{dB}$
Shadowing Loss (Satellite) $SF$ [32] 0  12 $0\text{ }-\text{ }12\text{ }\mathrm{dB}$
Line-of-Sight Probability (Terrestrial / Satellite) Refer to [31] / [32]
[…]
Urban/Rural UEs distribution proportion $40\%/60\%$
UE Antenna gain $G_{UE}$ [32] 0 $0\text{ }\mathrm{dBi}$
Satellite baseline energy consumption $E_{c}$ 500 $500\text{ }\mathrm{J}$
TABLE II: Simulation parameters.
V-B Optimization of the regularization parameter
As specified in Section III, $\lambda$ is a regularization parameter, which allows to control the trade-off between maximizing network SLT and minimizing the TN energy consumption (see (III)). Indeed, a high value of $\lambda$ leads BLASTER to increase the number of shutdown MBSs, as the priority becomes to reduce the TN energy consumption. Conversely, a low value of $\lambda$ indicates that the focus is on improving the SLT, by balancing the load and ensuring a proportional fair resource allocation. To achieve this goal, in our study, we set $\lambda$ inversely proportional to the number of UEs, $K$, in the network as follows:
$\lambda=\frac{\lambda_{\mathrm{max}}K_{\mathrm{min}}}{K},$ (28)
where $\lambda_{\mathrm{max}}$ is the value of $\lambda$ when the UEs number is equal to its lowest, i.e., $K_{\mathrm{min}}$. This approach allows us to simplify the setting of the hyper-parameter $\lambda$, by fixing $\lambda_{\rm max}$ and using (28). Therefore, in the following, we investigate the impact of $\lambda_{\rm max}$ on the behaviour of BLASTER. In particular, we let $\lambda_{\rm max}$ take values in $\left\{10,1e3,1e7,1e10\right\}$, and study the corresponding distribution of the SLT during high and low traffic as well as that of the total TN energy consumption.
In Fig. 4, we see a clear trend, with larger values of $\lambda_{\rm max}$ leading to a downward shift of the SLT distribution during high traffic. In fact, we observe a gradual increase of the median value of the SLT in high traffic, as $\lambda_{\rm max}$ decreases. This is in line with our assumption that a small value of $\lambda$ moves the focus from the TN energy consumption to improving the SLT. Specifically, in Fig. 4, we see a $2\,\%$ increase of SLT by varying $\lambda_{\rm max}$ from the largest value ($\lambda_{\rm max}=1e10$) to the smallest one ($\lambda_{\rm max}=10$). Note that, during high traffic, the performance in terms of SLT stagnates, once $\lambda_{\rm max}$ starts crossing extremely high values. Conversely, in Fig. 4, we notice that there is a limited change of the SLT, if the value of $\lambda_{\rm max}$ varies. In fact, in low traffic scenarios, our framework aims mainly to reduce the TN energy consumption, i.e., maximising the SLT has limited importance.
To verify this analysis, Fig. 5 plots the TN energy consumption for the distinct values of $\lambda_{\rm max}$, specified previously. We remark that the TN energy consumption increases throughout the day when the value of $\lambda_{\rm max}$ decreases. Indeed, a small value of $\lambda_{\rm max}$ hampers the transmit power optimization in (23), as each update becomes negligible, making it hard to effectively drive down the energy consumption of the TN.
Through these results, we are able to grasp the trade-off between reducing energy consumption and balancing the network load efficiently. Indeed, we observe that limiting the value of $\lambda_{\rm max}$, i.e., $\lambda_{\rm max}=10$ and $\lambda_{\rm max}=1e3$, we only achieve, during high traffic hours, a minor gain in terms of network SLT of approximately $2\%$ compared to the other cases ($\lambda_{\rm max}=1e7$ and $\lambda_{\rm max}=1e10$), while the total energy consumption surges by approximately $84\%$. Considering all of this, we fix $\lambda_{\rm max}=1e7$ as this value ensures a balanced performance in both SLT and TN energy consumption.
Figure 5: Daily profile of the terrestrial network energy consumption achieved by BLASTER for various $\lambda_{\rm max}$.
V-C Complexity and Convergence Analysis
In Fig. 6, we display the operational complexity of both proposed frameworks BLASTER and HEURISTIC as derived in Section IV-C. We notice that the operational complexity of BLASTER is higher than HEURISTIC throughout the day. This is expected, as the latter framework is implemented using less complex operations compared to the former, which resorts to multiple gradient descent methods. In fact, we see an average decrease in operational complexity of approximately $21\%$ in the day, which underlines the simplicity of HEURISTIC compared to BLASTER. Moreover, we notice that the complexity follows a pattern similar to the traffic load for both frameworks, which corroborates with the formulas derived in IV-C. Indeed, a lower number of total UEs $K$ naturally leads to a lower complexity, as seen in Fig. 6. On the opposite, a surge in traffic results in an increased complexity due to a higher number of operations needed to complete both frameworks.
To illustrate the convergence of BLASTER, we analyse the relative gain between successive iterations. A positive relative gain indicates an increase in the objective function, whereas a negative relative gain implies a decrease. Fig. 7 presents the average relative gain per iteration along with the corresponding standard deviation, computed across multiple runs of BLASTER at different hours of the day. Initially, we observe a sharp increase in relative gain, followed by a steady decline towards zero. Notably, the relative gain remains positive for the vast majority of iterations, confirming that the objective function generally improves over time. This behaviour highlights the convergence of BLASTER which on average occurs after approximately $30$ iterations.
Figure 6: Daily profile of the complexity of BLASTER and HEURISTIC.
Figure 7: Average relative gain per iteration for BLASTER.
V-D Impact of the Satellite Network on Traffic Distribution
In this section, we look at the shifting role that the satellites play on the mobile network throughout the day. Fig. 8 presents the daily profile of the fraction of UEs associated with the satellite network in the different frameworks under investigation. Also, we plot the hourly number of UEs deployed in the network, represented by the black dotted line. As we have previously underlined, the priority for the proposed BLASTER and HEURISTIC frameworks during low traffic is to reduce the TN energy consumption. Hence, the satellite becomes a compelling option for offloading the TN and shutting down lightly loaded MBSs. Fig. 8 shows that BLASTER and HEURISTIC achieve a $70\,\%$ and $500\,\%$ increase of the fraction of UEs associated to the satellite network in low-traffic hours as compared to the benchmark 3GPP-NTN. Also, we notice that the proportion of UEs associated to the satellite is far greater for HEURISTIC compared to BLASTER. This is due to the fact that HEURISTIC ensures that every UE that has a signal strength greater than $RSRP_{\rm min}$ is associated to the satellite, opposed to the more sophisticated BLASTER, who would still consider the available throughput before associating to the satellite. Conversely, during high traffic hours, we notice that the satellite network takes a less prominent role for both BLASTER and HEURISTIC, essentially acting as an umbrella, providing service to UEs that would otherwise be out of coverage. Accordingly, most of of the total bandwidth is allocated to the terrestrial tier, which can efficiently serve more UEs than the non-terrestrial tier and can significantly improve the spatial reuse.
Figure 8: Daily profile of the proportion of UEs associated to the satellite network.
V-E Analysis on the Network Sum Log-Throughput
In this section, we analyse the SLT achieved by the algorithms under investigation throughout the day and, in particular, how they adapt the network resources to the daily variations of the traffic load. Fig. 9 shows the daily profile of the SLT achieved by the schemes under investigation, and Fig. 10 presents the relative SLT gain for BLASTER, HEURISTIC and 3GPP-NTN compared to 3GPP-TN. Remember that in 3GPP-TN there are no satellites available to serve UEs perceiving low RSRP from terrestrial MBSs. In addition, the total bandwidth available for the TN is only $10$ MHz, which leads to low UE throughput, especially during busy hours. As expected, we observe an SLT performance improvement with the integration of satellites into the network. Indeed, LEO satellites are able to provide service to cell-edge UEs, which do not perceive a signal strong enough to be served from the terrestrial tier, thereby increasing the network SLT.
In more detail, in Fig. 9, we can see that both BLASTER and the proposed HEURISTIC outperform the 3GPP-NTN benchmark, with an average SLT increase of approximately $6\,\%$ across the day, highlighted in Fig. 10. Moreover, and although the SLT gain for 3GPP-NTN in Fig. 9 is more visually discernible during high-traffic hours, note that its relative gain with respect to 3GPP-TN, illustrated in Fig. 10 does not change drastically and remains around $2\%$ throughout the day.
In low traffic, Fig. 10 underlines that BLASTER outperforms HEURISTIC in terms of SLT. This is due to the different association methods used for both proposed frameworks. As detailed in Section V-D, the association criterion to the satellite in HEURISTIC is more lenient, which results to more UEs sharing the resources of the satellite, leading to a deteriorated throughput for its UEs and a worse SLT performance. Note that the improvement compared to the 3GPP benchmarks is more apparent during high traffic hours, typically midday to end of evening than low traffic hours. Indeed, as detailed before, the onus of both the HEURISTIC and BLASTER during low traffic is on reducing energy consumption, which explains the mitigated improvement of the network SLT.
During high traffic, both algorithms strive to balance the traffic load to maximize the SLT, which explains the striking improvement with respect to 3GPP-NTN and 3GPP-TN. That gain is due to various reasons. First of all, the UE-MBS association methods of HEURISTIC and BLASTER are designed to ensure a proportional fair resource allocation, which increases the SLT. In addition, the dynamic split of the network bandwidth based on the time-varying fraction of UEs associated to the satellite network (see Proposition 2) allows the proposed HEURISTIC and BLASTER to astutely distribute the frequency resources, differently than the 3GPP benchmarks [40]. In fact, since the majority of UEs are associated to a terrestrial MBS (see Fig. 8), a larger share of the resources are allocated to the terrestrial tier, which leads to a larger provided data-rate than the one achievable following the 3GPP benchmarks. By splitting the bandwidth based on the fraction of UEs associated to the satellites, we observe an enhancement of the mean SLT by at least $8\%$ and $6\%$ for both the proposed frameworks compared to 3GPP-TN and 3GPP-NTN respectively.
Figure 9: Daily profile of the sum log throughput.
Figure 10: Relative gain of the sum log throughput compared to 3GPP-TN.
V-F Analysis on the Network Energy Consumption
In this section, we study the performance of the proposed BLASTER and HEURISTIC in terms of energy usage. To this end, Fig. 11 displays the TN energy consumption throughout the day for the various algorithms under investigation. The red dotted line represents the energy consumption level for both 3GPP-TN and 3GPP-NTN settings, for which the TN energy consumption is at maximum level through the entire day, as they do not integrate neither power control nor MBS shutdown. Also, the black dotted line represents the same setting as 3GPP-TN, but with the added mechanism of shutting down all inactive MBSs when they have no UEs, denoted 3GPP--ENERGY SAVING. This setting reduces the average daily energy consumption by $17$ compared to 3GPP benchmarks, due to its ability to shut down MBSs, which leads to a more efficient energy use. As discussed in Section V-D, with BLASTER and HEURISTIC, the satellite network serves a larger proportion of UEs during low traffic than the standard 3GPP-NTN, which facilitates the shutdown of terrestrial MBSs. This is apparent in Fig. 11, as, in low traffic, the TN energy consumption sees an average decrease of approximately $67\,\%$ and $54\,\%$ for the BLASTER and HEURISTIC settings with respect to the benchmark 3GPP-NTN and 3GPP-TN, respectively. In comparison, for the 3GPP--ENERGY SAVING configuration, BLASTER and HEURISTIC experience a decrease of approximately $49\,\%$ and $31\,\%$, respectively. We also notice that, even though, during the low-traffic hours, the HEURISTIC is characterized by a larger share of UEs associated to the satellite network than the BLASTER, it still leads to a larger TN energy consumption. Interestingly, the transmit power optimization of BLASTER allows for a greater reduction of the transmit power of the TN MBSs. Indeed, the transmit power vector update step for BLASTER (21) - (26) is specifically designed to solve the optimization problem, reducing the terrestrial MBSs transmit power efficiently while the HEURISTIC updates the transmit power for terrestrial MBSs based on a rule of thumb, which suffices to reduce the network energy consumption. This also explains the striking difference in the high-traffic scenario, as the average energy consumed by the network is reduced by around $56\,\%$ and $14\,\%$ for BLASTER and HEURISTIC respectively, compared to the TN energy consumption in the 3GPP benchmarks. This result further underlines the flexible nature of both proposed frameworks, adjusting their behaviours to the traffic state and its specific demand. Moreover, Fig. 12 depicts the energy consumption of the satellite throughout the day for the settings detailed above. We notice that BLASTER and HEURISTIC considerably reduce the daily average energy consumption of the satellite by up to $80\%$ and $90\%$ respectively. It is important to note that the 3GPP benchmark is configured such that it utilizes the full $30$ MHz of bandwidth made available throughout the day, which leads to consistently higher energy consumption. In contrast, BLASTER and HEURISTIC use a smaller portion of the total bandwidth, as highlighted in Fig. 8 which results in lower energy usage. Additionally, by comparing Fig. 11 and Fig. 12, we notice that the energy consumption of a single satellite is negligible compared to the TN (less than $0.1\%$ of the total terrestrial energy use), which further validates our earlier assumption in Section III that the satellite energy consumption is insignificant in comparison to the TN.
Figure 11: Daily profile of the TN energy consumption.
Figure 12: Daily profile of the satellite energy consumption
VI Conclusions
In this paper, we have presented BLASTER, a framework designed to optimize radio resource management in an integrated TN-NTN. BLASTER aims to control UE association, splits the bandwidth between terrestrial and non-terrestrial tiers, and manages the MBS activation and MBS transmit power level. A novel method for splitting the bandwidth between terrestrial and non-terrestrial tiers is introduced, based on the fraction of UEs associated to the latter. The proposed algorithm also highlights the critical and dynamic part the satellites play in this integrated TN-NTN, adapting their role to various traffic demands. Indeed, the non-terrestrial tier takes a prominent role in low traffic, ensuring that terrestrial MBSs offload their UEs to the satellite to enable their shutdown. Conversely, the non-terrestrial tier embraces a secondary role in high traffic, mainly acting as an outlet to cell-edge UEs, facilitating load distribution while also giving up a share of its resources to the terrestrial tier. Simulation results display notable results, as the average TN energy consumption decreases by $67\,\%$ in low traffic compared to the 3GPP benchmarks, while the average SLT in high traffic increases by roughly $6\,\%$ compared to 3GPP-NTN. In future works, we plan to modify the current deterministic association method, as this method can be challenging to implement in practice. Instead, we aim to develop an ML-based solution, and adopt a distributed approach to tackle the problem. Another potential future enhancement could involve exploring a multi-beam system for the satellites, with a focus on studying inter-beam interference, and how to manage it efficiently, as well as considering the uplink (UL) scenario. Proof of Proposition 1. First, we inject (12) into formula (13) to get:
$\mathcal{D}\left(\mu\right)$ $=\mathcal{L}\left(X^{\star},\mu\right)$
$=\frac{1}{2}\|X^{\star}\|^{2}_{F}-\mathrm{Tr}\left({X^{\star}}^{T}\tilde{X}\right)+\frac{1}{2}\|\tilde{X}\|^{2}_{F}$
$\quad+\Big{[}\big{(}X^{\star}\odot\beta\big{)}\cdot p\Big{]}^{T}\mu-\left(\mathrm{RSRP}_{\mathrm{min}}\cdot\mathbbm{1}_{K}\right)^{T}\mu.$ (29)
Then, keeping the same notations as above, we notice that:
$\begin{split}\Big{[}\big{(}X^{\star}\odot\beta\big{)}\cdot p\Big{]}^{T}\mu=\mathrm{Tr}\left(X^{\star}\left(\beta\odot p^{\rm PAD}\odot\mu^{\rm PAD}\right)^{T}\right),\end{split}$ (30)
Indeed, developing the left component of (30), we get:
$\Big{[}\big{(}X^{\star}\odot\beta\big{)}\cdot p\Big{]}^{T}\mu=$
$\left[\begin{bmatrix}x_{11}^{*}&\cdots&x_{1L}^{*}\\ \vdots&\ddots&\vdots\\ x_{K1}^{*}&\cdots&x_{KL}^{*}\\ \end{bmatrix}\odot\begin{bmatrix}\beta_{11}&\cdots&\beta_{1L}\\ \vdots&\ddots&\vdots\\ \beta_{K1}&\cdots&\beta_{KL}\\ \end{bmatrix}\cdot\begin{bmatrix}p_{1}\\ \vdots\\ p_{L}\end{bmatrix}\right]^{T}\begin{bmatrix}\mu_{1}\\ \vdots\\ \mu_{L}\\ \end{bmatrix}$
$=\left[\begin{bmatrix}x_{11}^{*}\beta_{11}&\cdots&x_{1L}^{*}\beta_{1L}\\ \vdots&\ddots&\vdots\\ x_{K1}^{*}\beta_{K1}&\cdots&x_{KL}^{*}\beta_{KL}\\ \end{bmatrix}\cdot\begin{bmatrix}p_{1}\\ \vdots\\ p_{L}\end{bmatrix}\right]^{T}\begin{bmatrix}\mu_{1}\\ \vdots\\ \mu_{L}\\ \end{bmatrix}$
$=\begin{bmatrix}\sum\limits_{j=1}^{L}x^{*}_{1j}\beta_{1j}p_{j}\\ \vdots\\ \sum\limits_{j=1}^{L}x^{*}_{Kj}\beta_{Kj}p_{j}\end{bmatrix}^{T}\begin{bmatrix}\mu_{1}\\ \vdots\\ \mu_{L}\\ \end{bmatrix}=\sum\limits_{i=1}^{K}\mu_{i}\left(\sum\limits_{j=1}^{L}x^{*}_{ij}\beta_{ij}p_{j}\right)$
Then, focusing on the right component of (30), we have:
$X^{\star}\left(\beta\odot p^{\rm PAD}\odot\mu^{\rm PAD}\right)^{T}=\begin{bmatrix}x_{11}^{*}&\cdots&x_{1L}^{*}\\ \vdots&\ddots&\vdots\\ x_{K1}^{*}&\cdots&x_{KL}^{*}\\ \end{bmatrix}\cdot$
$\left(\begin{bmatrix}\beta_{11}&\cdots&\beta_{1L}\\ \vdots&\ddots&\vdots\\ \beta_{K1}&\cdots&\beta_{KL}\\ \end{bmatrix}\odot\begin{bmatrix}p_{1}&\cdots&p_{L}\\ \vdots&\ddots&\vdots\\ p_{1}&\cdots&p_{L}\\ \end{bmatrix}\odot\begin{bmatrix}\mu_{1}&\cdots&\mu_{1}\\ \vdots&\ddots&\vdots\\ \mu_{K}&\cdots&\mu_{K}\\ \end{bmatrix}\right)^{T}$
$=\begin{bmatrix}x_{11}^{*}&\cdots&x_{1L}^{*}\\ \vdots&\ddots&\vdots\\ x_{K1}^{*}&\cdots&x_{KL}^{*}\\ \end{bmatrix}\begin{bmatrix}\beta_{11}p_{1}\mu_{1}&\cdots&\beta_{1L}p_{L}\mu_{1}\\ \vdots&\ddots&\vdots\\ \beta_{K1}p_{1}\mu_{K}&\cdots&\beta_{KL}p_{L}\mu_{K}\\ \end{bmatrix}^{T}$
$=\begin{bmatrix}x_{11}^{*}&\cdots&x_{1L}^{*}\\ \vdots&\ddots&\vdots\\ x_{K1}^{*}&\cdots&x_{KL}^{*}\\ \end{bmatrix}\begin{bmatrix}\beta_{11}p_{1}\mu_{1}&\cdots&\beta_{K1}p_{1}\mu_{K}\\ \vdots&\ddots&\vdots\\ \beta_{1L}p_{L}\mu_{1}&\cdots&\beta_{KL}p_{L}\mu_{K}\\ \end{bmatrix}$
$=\begin{bmatrix}\sum\limits_{j=1}^{L}x^{*}_{1j}\left(\beta_{1j}p_{j}\mu_{1}\right)&\cdots&\cdots\\ \vdots&\ddots&\vdots\\ \cdots&\cdots&\sum\limits_{j=1}^{L}x^{*}_{Kj}\left(\beta_{Kj}p_{j}\mu_{K}\right)\\ \end{bmatrix}$
Then, by applying the trace operator, we effectively get:
$\mathrm{Tr}\left(X^{\star}\left(\beta\odot p^{\rm PAD}\odot\mu^{\rm PAD}\right)^{T}\right)=\sum\limits_{i=1}^{K}\mu_{i}\left(\sum\limits_{j=1}^{L}x^{*}_{ij}\beta_{ij}p_{j}\right)$
Injecting (30) into (VI) and noticing that , we obtain:
$\begin{split}\mathcal{D}\left(\mu\right)&=\frac{1}{2}||X^{\star}||^{2}_{F}-\mathrm{Tr}\left(X^{\star}\left[\tilde{X}-\beta\odot p^{\rm PAD}\odot\mu^{\rm PAD}\right]^{T}\right)\\ &-\left(RSRP_{\mathrm{min}}\cdot\mathbbm{1}_{K}\right)^{T}\mu,\end{split}$ (31)
which concludes the proof of the proposition.